OpenStudy (anonymous):

Find the inverse. g(x) = 2x + 5 Please help! Studying for a test

OpenStudy (anonymous):

\[g^-1(x)=\frac{- 5 }{ 2 }+\frac{ x }{ 2 }\]

OpenStudy (anonymous):

The steps on doing it? I will thank you if you walked me through them lol

OpenStudy (anonymous):

g(x)= 2x+5 (solve for "x=") g(x) -5 = 2x \[\frac{ g(x)-5 }{ 2 }=x\] (switch the x for the y (g(x)) \[y=\frac{ x-5 }{ 2 }\]

OpenStudy (anonymous):

was this what you were looking for?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

No problem!